# infomeasure: A Comprehensive Python Package for Information Theory Measures and Estimators

### v1

@@ Introduction
Information theory, i.e. the mathematical study of information transmission and processing [1, 2, 3], has become a cornerstone of modern data analysis, with applications ranging from the study of molecules [4, 5] to outer space [6]. It has further found natural applicability in the study of complex systems [7], i.e. those systems composed of numerous interacting components; any attempt to characterise their dynamics must necessarily involve information theory, as the latter is the language of computation, and this defines their nature [8]. Metrics can then be used to describe interactions and information flows between these elements, and, more generally, internal organisation patterns. While the prototypical example of this may be neuroscience [9, 10, 11], relevant applications can be found in fields as diverse as finance [12] or ecology [13].
These examples may give the impression that applying information theory measures to a given problem is a simple process; this is nevertheless not the case, with more widespread applications being hindered by four main barriers. Firstly, the measurement of a given information aspect, e.g. uncertainty or information transmission, relies on the use of estimators. These are functions and algorithms that yield an estimation of the probability distributions from which data are generated, using a limited number of observed samples. Notably, limited theoretical guidelines are available to select the best estimator for a given real problem; using suboptimal ones may result in biases and wrong results. Still, this is made more challenging by the second barrier: while multiple software libraries exist, they usually focus on specific measures and/or estimators. This, together with the fact that many of them are not constantly maintained (see Tab. 1 for an overview), makes any comparison (e.g. testing multiple estimators) more challenging. The consequence is that a practitioner may only be able to compare a few estimators at best, unless multiple toolboxes are integrated in the processing pipeline. Thirdly, the large variety of available software tools is an obstacle towards reproducibility, as they may include variations in specific aspects of the implementation that may affect the obtained results. Lastly, the practitioner will face the problem of the large computational cost of these measures; unless robust and highly optimised algorithms are available, the application to large-scale data sets is challenging at best.
@@
Table 1: Overview of existing information theory toolboxes. The last column reports the date of the last human commit (i.e. update), as of May 2025.
@@
dit https://dit.readthedocs.io/en/latest/index.html Python Discrete April 2025
infotheory https://mcandadai.com/infotheory/ Python/C++ Discrete/continuous, limited estimators. August 2020
InfoTheory.jl https://github.com/robertfeldt/InfoTheory.jl Julia Entropy only January 2016
JIDT https://jlizier.github.io/jidt/ Java Discrete / Continuous May 2025
pyentropy https://code.google.com/archive/p/pyentropy/ Python Entropy and MI October 2019
pyEntropy (pyentrp) https://github.com/nikdon/pyEntropy Python Entropy July 2024
PyInform https://pypi.org/project/pyinform/ Python Discrete December 2019
Pyitlib (MIT) https://pypi.org/project/pyitlib/ Python Discrete July 2023
The Transfer Entropy Toolbox (TET) https://code.google.com/archive/p/transfer-entropy-toolbox/ Matlab Binary time series June 2013
TRENTOL https://trentool.github.io/TRENTOOL3/ Matlab Discrete/continuous, only TE, limited estimators November 2017
@@
  • Joint Entropy. Amount of information gained when jointly observing two random variables $X$ and $Y$.
@@
As previously introduced, these measures can only be calculated by knowing the probability distributions underlying the observed data. As these are usually not known, except for the case of simple toy models, the solution requires the use of estimators, i.e. functions and algorithms that yield an estimation of the probability distributions from which data are generated. The infomeasure library supports a large set of them, whose main families are listed and described below. Additionally, most combinations of estimators and measures are natively supported, see Tab. 2.
@@
Table 2: Information Measures and Estimation Methods. Measures marked with an asterisk (*) can also be computed using Rényi and Tsallis entropy formulations, in addition to the standard Shannon entropy. The non-standard cross-entropy notation has been borrowed by from Christopher Olah’s blog post “Visual Information Theory” [24]. We choose this nomenclature $H_{Q}(P)$, as the widely used $H(P,Q)$ is ambiguous with joint entropy.
Measures Notation Discrete Kernel Metric / kNN Ordinal
Estimator Estimator Estimator Estimator Estimator
@@
Kullback-Leibler Divergence (KLD) $\operatorname{KLD}(P||Q)$
Jensen-Shannon Divergence (JSD) $\operatorname{JSD}(P||Q)$
Some additional features of the library ought to be here highlighted. From the viewpoint of usability, the package unifies the slicing mechanism for the TE estimators, ensuring consistency across all approaches and minimising potential points of failure. It can be used to obtain local values for the entropy, MI, and TE; thus allowing to describe the evolution of the information transfer through time [25]. It further simplifies hypothesis testing, by yielding $p$-values and $t$-scores for MI and TE. Finally, the TE implements an Effective TE (eTE) variant, reducing biases from finite sample sizes [26].
@@
Figure 1: Evolution of the computational cost as a function of the time series length. Results have been obtained with CPython 3.13.2 and Clang 18.1.8 (Darwin 24.4.0) on an Apple M4 Pro; only one core used in the computations. Points and whiskers respectively indicate the average and standard deviation.
@@
data = [0, 1, 0, 1, 0, 1, 0, 1]
entropy = im.entropy(data, approach="kernel", bandwidth=3, kernel="box")
@@
We start with the simplest case of a random variable $X$ defined by a normal distribution of zero average and $\sigma$ standard deviation; its maximum likelihood entropy is known to be equal to $H(X)=1/2\log(2\pi e\sigma^{2})$. The top left panel of Fig. 2 reports the evolution of such entropy as a function of $\sigma$ (black line), and of the results obtained by different estimators. All methods yield very good approximations of the analytical values, the only exceptions being the ordinal and discrete estimators. In the case of the former, the error comes from the discretisation of values to the nearest integer, which destroys all information for small values of $\sigma$. On the other hand, the ordinal symbolisation process only considers the relative amplitude of values within a pattern, but not their absolute amplitude; in other words, the obtained entropy is independent of the scaling introduced by $\sigma$, hence the constant value.
We next move to a bivariate case, by evaluating the MI between two Gaussian random variables $X$ and $Y$, both with zero mean and unit variance, and with a linear correlation coefficient between them of $\rho$. The theory indicates that the MI between them is given by $I(X;Y)=-1/2\log(1-\rho^{2})$. Again, results obtained by the estimators match the theoretical prediction, see top right panel of Fig. 2. In this case the exception is given by Tsallis’ MI, here due to the use of $q=1.05$–note that, for $q=1$, the Tsallis’ MI becomes identical to the Shannon’s one.
@@
The interested reader can find the code to reproduce these four results, as well as some additional considerations, in the documentation of the package, see https://infomeasure.readthedocs.io/en/0.4.0/demos/gaussian_data/ and https://infomeasure.readthedocs.io/en/0.4.0/demos/Schreiber_Article/.
@@
As a final topic and example of use case, we here describe the task of analysing information transmission in human brain time series. The underlying idea, at the foundation of the so-called “network neuroscience” field [27, 28], is that brain activity can be represented by networks; their nodes then map different regions of the brain, and a pair of them are connected by links whenever a statistical dependency is observed between the corresponding dynamics [29, 30, 31]. We here reconstruct functional networks by starting from electroencephalographic (EEG) recordings of 14 patients suffering from schizophrenia, and of the same number of matched control subjects [32], freely available at http://dx.doi.org/10.18150/repod.0107441. Data correspond to a resting state, i.e. with the subjects doing no cognitive tasks and with their eyes closed; 19 electrodes were recorded at 250Hz, following the standard positions: Fp1, Fp2, F7, F3, Fz, F4, F8, T3, C3, Cz, C4, T4, T5, P3, Pz, P4, T6, O1, O2. Time series have not been preprocessed and have been used as provided [32].
We start by comparing the results obtained by calculating the Transfer Entropy measure between pairs of time series, using three different estimators: the box kernel, the KSG, and the ordinal one. We specifically analyse the average information detected as a function of the lag $\tau$, for different parameters and across $100$ random pairs of time series. Results are reported in the left panels of Fig. 3; for the sake of clarity, the TE has been normalised according to the largest value detected by each estimator. Several interesting ideas can already be drawn. Firstly, changing the parameter of each estimator has a varying effect: it strongly modifies the results in the box kernel and ordinal cases, but has no impact in the KSG case. Secondly, in all cases two maxima are observed, one for small $\tau$s, and a second one around $\tau\approx 15$. While the former may be the result of the smoothness of the time series, the latter probably represents a more genuine information transfer. We then choose the parameters yielding the clearest peak: $\tau=17$ and $s=0.1$ for the box kernel; $\tau=17$ and $k=4$ for KSG; and $\tau=15$ and $D=4$ for the ordinal estimator.
@@
Panels in the rightmost column finally report the differences that each estimator observes between control subjects and patients. Similarly to the previous case, green (respectively, red) shades indicate a higher (lower) TE in the control subjects. Estimators here have completely different behaviours: the box kernel detects differences, but no clear patterns in them; the KSG, minimal differences between the two groups; and the ordinal estimator, a substantially higher TE for the control subjects across all regions, especially marked in the central lobe. Note that this latter result is aligned with previous findings in the literature [33, 34, 35].
@@
In short, this use case illustrates some of the advantages of the infomeasure library. On the one hand, it simplifies the computation of information-theoretical measures, in the limit that only one line of code had to be changed to alternate approach and parameters to obtain all results presented in Fig. 3. On the other hand, such flexibility enables comparisons beyond what is usually found in the literature: as an example, we here showed that different estimators observe different patterns discriminating between control subjects and schizophrenic patients.
@@
We here presented infomeasure, a comprehensive Python library for the efficient and accurate computation of information-theoretic measures. Compared to alternative libraries, it presents multiple advantages, including: the availability of numerous estimators, and the possibility of plug them in the calculation of numerous measures (see Tab. 2); a modular structure that adapts to the expertise of the user, allowing to get results from one line of code, or by delving deeper into individual modules; the possibility of obtaining local values for the measures; and the support of hypothesis testing, through $p$-values and $t$-scores for MI and TE.
@@ Acknowledgements
Accession codes: The infomeasure package (version 0.4.0) is archived on Zenodo, at https://doi.org/10.5281/zenodo.15330677.

### v2

@@ Introduction
Information theory, i.e. the mathematical study of information transmission and processing [1, 2, 3], has become a cornerstone of modern data analysis, with applications ranging from the study of molecules [4, 5] to outer space [6]. It has further found natural applicability in the study of complex systems [7], i.e. those systems composed of numerous interacting components; any attempt to characterise their dynamics often involves information theory, as the latter is the language of computation, and this defines their nature [8]. Metrics can then be used to describe interactions and information flows between these elements, and, more generally, internal organisation patterns. While the prototypical example of this may be neuroscience [9, 10, 11], relevant applications can be found in fields as diverse as finance [12], ecology [13], or machine [14] and deep learning [15].
These examples may give the impression that applying information theory measures to a given problem is a simple process; this is nevertheless not the case, with more widespread applications being hindered by four main barriers. Firstly, the measurement of a given information aspect, e.g. uncertainty or information transmission, relies on the use of estimators. These are functions and algorithms that yield an estimation of the probability distributions from which data are generated, using a limited number of observed samples. Specifically, limited theoretical guidelines are available to select the best estimator for a given real problem; using suboptimal ones may result in biases and wrong results. Still, this is made more challenging by the second barrier: while multiple software libraries exist, they usually focus on specific measures and/or estimators. The interested reader will find a list of some toolboxes, along with their key characteristics, in Tab. 1; while it is not exhaustive, and does not include application-specific libraries (e.g. tailored to neuroscience data), it clearly shows the heterogeneity both in metrics and in update frequencies. This variety makes any comparison (e.g. testing multiple estimators) more challenging. The consequence is that a practitioner may only be able to compare a few estimators at best, unless multiple toolboxes are integrated in the processing pipeline. Thirdly, the large variety of available software tools is an obstacle towards reproducibility, as they may include variations in specific aspects of the implementation that may affect the obtained results. Lastly, the practitioner will face the problem of the large computational cost of these measures; unless robust and highly optimised algorithms are available, the application to large-scale data sets is challenging at best.
@@
Table 1: Overview of existing information theory toolboxes. This list is non-exhaustive. The last column reports the date of the last human commit on the main branch, as of July 2025.
@@
dit[16] https://dit.readthedocs.io/en/latest/index.html Python Discrete May 2025
DiscreteEntropy.jl[17] https://kellino.github.io/DiscreteEntropy.jl/dev/ Julia Discrete May 2025
IDTxl[18] https://github.com/pwollstadt/IDTxl Python MI, TE, AIS, PID April 2025
infotheory[19] https://mcandadai.com/infotheory/ Python/C++ Discrete/continuous, limited estimators. August 2020
InfoTheory.jl[20] https://github.com/robertfeldt/InfoTheory.jl Julia Entropy only January 2016
JIDT[21] https://jlizier.github.io/jidt/ Java Discrete / Continuous May 2025
pyentropy[22] https://code.google.com/archive/p/pyentropy/ Python Entropy and MI October 2019
pyEntropy (pyentrp)[23] https://github.com/nikdon/pyEntropy Python Entropy July 2025
PyInform[24] https://pypi.org/project/pyinform/ Python Discrete December 2019
Pyitlib (MIT)[25] https://pypi.org/project/pyitlib/ Python Discrete May 2025
The Transfer Entropy Toolbox (TET)[26] https://code.google.com/archive/p/transfer-entropy-toolbox/ Matlab Binary time series June 2013
TRENTOL[27] https://trentool.github.io/TRENTOOL3/ Matlab Discrete/continuous, only TE, limited estimators November 2017
@@
  • Joint Entropy. Amount of information needed to describe two random variables $X$ and $Y$ together, or equivalently, a measure of the total uncertainty in their joint outcomes.
@@
As previously introduced, these measures can only be calculated by knowing the probability distributions underlying the observed data. As these are usually not known, except for special cases and toy models, the solution requires the use of estimators, i.e. functions and algorithms that yield an estimation of the probability distributions from which data are generated. The infomeasure library supports a large set of them, whose main families are listed and described below. Additionally, most combinations of estimators and measures are natively supported, see Tab. 2. Finally, a short guide for the user new to these approaches is provided at https://infomeasure.readthedocs.io/en/latest/guide/estimator_selection/.
@@
  • Bias-corrected estimators. Set of estimators substituting the plug-in one for the Shannon’s entropy, designed to provide a lower bias for small sample sizes. These estimators have been selected based on the review included in Refs. [38, 39].
Table 2: Information Measures and Estimation Methods. Measures marked with an asterisk (*) can also be computed using Rényi and Tsallis entropy formulations, in addition to the standard Shannon entropy. The non-standard cross-entropy notation has been borrowed from Christopher Olah’s blog post “Visual Information Theory” [40]. We choose this nomenclature $H_{Q}(P)$, as the widely used $H(P,Q)$ is ambiguous with joint entropy. For the bias-corrected estimators, measures marked with a dagger ($\dagger$) are only available for a subset of the estimators.
Measures Notation Discrete Kernel Metric / kNN Ordinal Bias-corrected
Estimator Estimator Estimator Estimator Estimator Estimators
@@
Kullback-Leibler Divergence (KLD)* $\operatorname{KLD}(P||Q)$ $\dagger$
Jensen-Shannon Divergence (JSD) $\operatorname{JSD}(P||Q)$ $\dagger$
Some additional features of the library ought to be here highlighted. From the viewpoint of usability, the package unifies the slicing mechanism for the TE estimators, ensuring consistency across all approaches and minimising potential points of failure. It can be used to obtain local values for the entropy, MI, and TE; thus allowing to describe the evolution of the information transfer through time [41]. It further simplifies hypothesis testing, by yielding $p$-values, $t$-scores and confidence intervals for MI and TE. Finally, the TE implements an Effective TE (eTE) variant, reducing biases from finite sample sizes [42].
@@
Figure 1: Evolution of the computational cost as a function of the time series length. Results have been obtained with CPython 3.13.2 and Clang 18.1.8 (Darwin 24.4.0) on an Apple M4 Pro; only one core used in the computations. Points and whiskers respectively indicate the average and standard deviation over ten independent realisations. Equivalent results for bias-corrected estimators are available at https://infomeasure.readthedocs.io/en/latest/demos/Time_Performance/.
@@
data = [1.24, 0.92, 1.87, 1.51, 0.48, 3.60, 1.32]
entropy = im.entropy(data, approach="kernel", bandwidth=2, kernel="box")
@@
We start with the simplest case of a random variable $X$ defined by a normal distribution of zero average and $\sigma$ standard deviation; its maximum likelihood entropy is known to be equal to $H(X)=1/2\log(2\pi e\sigma^{2})$. The top left panel of Fig. 2 reports the evolution of such entropy as a function of $\sigma$ (black line), and of the results obtained by different estimators. All methods yield very good approximations of the analytical values, the only exceptions being the ordinal and discrete estimators. In the case of the former, the error comes from the discretisation of values to the nearest integer, which destroys all information for small values of $\sigma$. On the other hand, the ordinal symbolisation process only considers the relative amplitude of values within a pattern, but not their absolute amplitude; in other words, the obtained entropy is independent of the scaling introduced by $\sigma$, hence the constant value. Finally, the box kernel estimator deviates for very small standard deviations, resulting from the kernel bandwidth being too small to capture the information; yet it recovers the analytical curve for larger values.
We next move to a bivariate case, by evaluating the MI between two Gaussian random variables $X$ and $Y$, both with zero mean and unit variance, and with a linear correlation coefficient between them of $\rho$. The theory indicates that the MI between them is given by $I(X;Y)=-1/2\log(1-\rho^{2})$. Again, results obtained by the estimators match the theoretical prediction, see top right panel of Fig. 2. In this case the exception is given by Tsallis’ MI, here due to the use of q=1.05q=1.05–note that, for $q=1$, the Tsallis’ MI becomes identical to the Shannon’s one.
@@
The interested reader can find the code to reproduce these four results, as well as some additional considerations, in the documentation of the package, see https://infomeasure.readthedocs.io/en/0.5.0/demos/gaussian_data/ and https://infomeasure.readthedocs.io/en/0.5.0/demos/Schreiber_Article/.
@@
As a final topic and example of use case, we here describe the task of analysing information transmission in human brain time series. This section is independent from the remainder of the text; the reader not interested in it can safely skip to the conclusions. The underlying idea, at the foundation of the so-called “network neuroscience” field [43, 44], is that brain activity can be represented by networks; their nodes then map different regions of the brain, and a pair of them are connected by links whenever a statistical dependency is observed between the corresponding dynamics [45, 46, 47]. We here reconstruct functional networks by starting from electroencephalographic (EEG) recordings of 14 patients suffering from schizophrenia, and of the same number of matched control subjects [48], freely available at http://dx.doi.org/10.18150/repod.0107441. Data correspond to a resting state, i.e. with the subjects doing no cognitive tasks and with their eyes closed; 19 electrodes were recorded at 250Hz, following the standard positions: Fp1, Fp2, F7, F3, Fz, F4, F8, T3, C3, Cz, C4, T4, T5, P3, Pz, P4, T6, O1, O2. Time series have not been preprocessed and have been used as provided [48].
[…]
@@
Panels in the rightmost column finally report the differences that each estimator observes between control subjects and patients. Given the average value observed in each group and for each pair of electrodes $i$ and $j$, the difference has been estimated as $\delta_{i,j}=\log_{2}\overline{TE_{i,j}^{cntr}}/\overline{TE_{i,j}^{sch}}$. Positive values of $\delta_{i,j}$ thus indicate that the TE is on average higher in control subjects (represented as green squares); conversely, negative values (red squares) indicate a higher value for patients. Estimators here have completely different behaviours: the box kernel detects differences, but no clear patterns in them; the KSG, minimal differences between the two groups; and the ordinal estimator, a substantially higher TE for the control subjects across all regions, especially marked in the central lobe. Note that this latter result is aligned with previous findings in the literature [49, 50, 51].
@@
In short, this use case illustrates some of the advantages of the infomeasure library. On the one hand, it simplifies the computation of information-theoretical measures, in the limit that only one line of code had to be changed to alternate approach and parameters to obtain all results presented in Fig. 3. On the other hand, such flexibility enables comparisons beyond what is usually found in the literature: as an example, we here showed that different estimators observe different patterns discriminating between control subjects and schizophrenic patients. In any case, the reader should note that this is a basic analysis that does not substitute more detailed studies, which would require better data pre-processing and artefact elimination, statistical analysis of the differences, and so forth [49, 50, 51].
@@
We here presented infomeasure, a comprehensive Python library for the efficient and accurate computation of information-theoretic measures. Compared to alternative libraries, it presents multiple advantages, including: the availability of numerous estimators, and the possibility of plug them in the calculation of numerous measures (see Tab. 2); a modular structure that adapts to the expertise of the user, allowing to get results from one line of code, or by delving deeper into individual modules; the possibility of obtaining local values for the measures; and the support of hypothesis testing, through $p$-values and $t$-scores for MI and TE. While some elements are not automatised and are left to the control of the user (as e.g. the selection of the best parameters), the structure of the library simplifies its integration into larger data analysis pipelines. Additionally, we plan to continuously expand the set of metrics and estimators - the interested reader can find a future development roadmap in the documentation.
@@ Acknowledgements
Accession codes: The infomeasure package (version 0.5.0) is archived on Zenodo, at https://zenodo.org/records/15792893.
@@
Data availability
The dataset analysed during the current study is available in the RepOD repository, http://dx.doi.org/10.18150/repod.0107441.